OpenStudy (anonymous):

6 over 16 = 30 over ?

mathslover (mathslover):

\[6/16 = 30/? \] let ? be x therefore we have \[6/16=30/x\] \[6x=30*16\] \[x=30*16/6\] \[x=5*16\] \[x=80\]

OpenStudy (anonymous):

what do you multiply 6 by to get 30?

Parth (parthkohli):

@satellite73 Why do we even want that number?

Parth (parthkohli):

Cross multiply

OpenStudy (anonymous):

whatever that number is, multiply 16 by it to get your answer

Parth (parthkohli):

lol what a cover up